# VLA-JEPA: Enhancing Vision-Language-Action Model with Latent World Model

### v1

@@ 1 Introduction
  • VLA-JEPA: leakage-free, state-level JEPA pretraining for VLA. We propose a JEPA-style latent predictive alignment scheme that learns action-relevant transition semantics by predicting and aligning future latent states—without pixel reconstruction, information leakage and only one-stage pretraining pipeline.
@@ 2.1 Vision-Language-Action Models
With the rapid advancement of Large Language Models (LLMs) [llava23, karamcheti2024prismatic, beyer2024paligemma, DreamLLM23, yang2025qwen3, ShapeLLM24] and large-scale robot datasets [o2023open, ebert2021bridge, khazatsky2024droid, deng2025graspvla], Vision-Language-Action (VLA) models have become a dominant paradigm in robot learning [kim2024openvla, 24pi0, xu2025seeing]. The RT series [RT123, RT223, RTH24] pioneered fine-tuning multimodal LLMs on robot demonstrations, and subsequent works further improved manipulation and navigation performance [kim2024openvla, 24pi0, qu25spatialvla, liang2025discrete, zhang2024navid, zhang2024uni, pretrainingdisentangled, wu2026pragmatic]. However, most VLA methods rely heavily on large-scale action-labeled robot data, which is costly and difficult to scale [wang2025vla]. To reduce dependence on explicit action supervision, recent works introduce multimodal Chain-of-Thought signals, such as hierarchical planning [RTH24, lin2025onetwovla, sofar25, zhou2025chatvla, ji2025robobrain], subgoal or rollout prediction [seer24, cotvla25, 3dvla, wang2025unified, lin2025hif, chen2025unified, zhang2025future, liu2025mla], object-centric conditioning [deng2025graspvla, 25pi0.5, ranasinghe2025pixel, zhong2025flowvla, upvla25, lv2025f1, cai2026internvla] and latent future embeddings or actions [go125bu, dywa25, zhang2025dreamvla, liu2026last, zhong2026acot]. Nevertheless, these approaches suffer from relying on action-labeled data. In contrast, VLA-JEPA learns action-centric representations via latent predictive alignment, avoiding explicit future reconstruction and reducing the need for large-scale action supervision.
@@ 2.2 Latent action learning for robotics
To leverage large-scale videos without action labels, ILPO [edwards2019imitating], LAPO [23lapo] and Genie [bruce2024genie] propose using latent action for video games. For robot learning, LAPA [ye2024lapa], IGOR [chen2024igor], UniVLA [bu2025univla], MotoGPT [moto24], Adaworld [gaoa24daworld], CoMo [yang2025como] and StaMo [liu2025stamo] extract discrete or continuous motion tokens from frame transitions, and pretrain VLA to predict these latent actions before mapping them to real robot controls. To align the latent action to the real action space, villa-x [chen2025villa], XR-1 [fan2025xr], CLAP [zhang2026clapcontrastivelatentaction] and VITA [ma2025unifying] propose that extract latent action from both robot and human video and use a unified codebook. However, since latent actions are often learned directly from adjacent frames, models may exploit pixel-level shortcuts and encode future-frame leakage. Although LAOF [bu2025laof] and Motus [bi2025motus] propose to use optical flow to constrain the space of latent action, it still entangles controllable dynamics with camera motion and background changes. As a result, learned latent spaces may align more with visual deltas than with actionable control signals, requiring multi-stage training pipelines and additional alignment mechanisms [zhang2025latent]. On the contrary, our proposed VLA-JEPA learns action-relevant representations without relying on delta frame info extraction, avoiding leakage and pixel shortcuts while enabling single-stage end-to-end pretraining.
@@ 4 Experiments
Spatial Object Goal 10 Avg
@@ 4.5 Further Analysis and Ablation Study
Q3: How does the number of future video horizon affect performance?

### v2

@@ 1 Introduction
  • VLA-JEPA: leakage-free, state-level JEPA pretraining. We propose a JEPA-style latent predictive alignment scheme that learns action-relevant transition semantics by predicting and aligning future latent states—without pixel reconstruction, information leakage and only one-stage pretraining pipeline.
@@ 2.1 Vision-Language-Action Models
With the rapid advancement of Large Language Models (LLMs) [llava23, karamcheti2024prismatic, beyer2024paligemma, DreamLLM23, yang2025qwen3, ShapeLLM24] and large-scale robot datasets [o2023open, ebert2021bridge, khazatsky2024droid, deng2025graspvla], Vision-Language-Action (VLA) models have become a dominant paradigm in robot learning [kim2024openvla, 24pi0, xu2025seeing, Chen2025SimpleVLA-RL]. The RT series [RT123, RT223, RTH24] pioneered fine-tuning multimodal LLMs on robot demonstrations, and subsequent works further improved manipulation and navigation performance [kim2024openvla, 24pi0, qu25spatialvla, liang2025discrete, zhang2024navid, zhang2024uni, pretrainingdisentangled, wu2026pragmatic]. However, most VLA methods rely heavily on large-scale action-labeled robot data, which is costly and difficult to scale [wang2025vla]. To reduce dependence on explicit action supervision, recent works introduce multimodal Chain-of-Thought signals, such as hierarchical planning [RTH24, lin2025onetwovla, sofar25, zhou2025chatvla, ji2025robobrain], subgoal or rollout prediction [seer24, cotvla25, 3dvla, wang2025unified, lin2025hif, chen2025unified, zhang2025future, liu2025mla], object-centric conditioning [deng2025graspvla, 25pi0.5, ranasinghe2025pixel, zhong2025flowvla, upvla25, lv2025f1, cai2026internvla] and latent future embeddings or actions [go125bu, dywa25, zhang2025dreamvla, liu2026last, zhong2026acot]. Nevertheless, these approaches suffer from relying on action-labeled data. In contrast, VLA-JEPA learns action-centric representations via latent predictive alignment, avoiding explicit future reconstruction and reducing the need for large-scale action supervision.
@@ 2.2 Latent action learning for robotics
To leverage large-scale videos without action labels, ILPO [edwards2019imitating], LAPO [23lapo] and Genie [bruce2024genie] propose using latent action for video games. For robot learning, LAPA [ye2024lapa], IGOR [chen2024igor], UniVLA [bu2025univla], MotoGPT [moto24], Adaworld [gaoa24daworld], CoMo [yang2025como] and StaMo [liu2025stamo] extract discrete or continuous motion tokens from frame transitions, and pretrain VLA to predict these latent actions before mapping them to real robot controls. To align the latent action to the real action space, villa-x [chen2025villa], XR-1 [fan2025xr], CLAP [zhang2026clapcontrastivelatentaction] and VITA [ma2025unifying] propose that extract latent action from both robot and human video and use a unified codebook. However, since latent actions are often learned directly from adjacent frames, models may exploit pixel-level shortcuts and encode future-frame leakage. Although some methods attempt to mitigate the above issues by constraining the latent action space—through optical-flow [bu2025laof, bi2025motus], or object-centric constraints [Alexander2025Object], among others [Alexander2025LALRS, Alexander2026VLM]—such constraints inevitably bias the learned latent actions toward aligning with hand-crafted, visually defined priors rather than the true underlying controllable factors. Consequently, when such priors become mismatched in novel environments, the learned latent actions tend to fail systematically, and the latent spaces may align more with visual deltas than with actionable control signals, necessitating multi-stage training pipelines and additional alignment mechanisms [zhang2025latent]. On the contrary, our proposed VLA-JEPA learns action-relevant representations without relying on delta frame info extraction, avoiding leakage and pixel shortcuts while enabling single-stage end-to-end pretraining.
@@ 4 Experiments
Spatial Object Goal LIBERO-10 Avg
@@ 4.5 Further Analysis and Ablation Study
Q3: How does the number of future video horizon affect performance?